If one leg of a right triangle measures 3 centimeters and the other leg measures 6 centimeters, what is the legnth of the hypotenuse? This is a multiple choice A. 3 to the square root of 5 centimeters B. 5 to the square root of 3 centimeters C. 9 centimeters D. the square root of 5 centimeters? Please help me, thank you very much!
Answer by arunpaul(104)

let the three side of triangle be A,B,C
and C is the hypotenuse and A=3 and B =6;
so by the law of right angled triangle
A^2+B^2=C^2
so C = SQRT(A^2+B^2) (where ^ indicate power)
C = SQRT (3^2+6^2)
=SQRT(9+36)
=SQRT(45)
=SQRT(9*5) ( where * mean multiplication)
=3square root of 5 (ANS)
so the answer is the choice A